Circular No. 7264

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1999 S3

M. Bezpalko reports the discovery by LINEAR of a comet. Additional observations have been reported following posting on the NEO Confirmation Page:

1999 UT  α_{2000}  δ_{2000}  m_{1} Observer
Sept.24.34373 1^{h}21^{m}56.^{s}90 +46^{o}35′31′′.1  Bezpalko
24.35812 1 21 53.29 +46 36 10.2  _{”}
24.38689 1 21 46.26 +46 37 29.8  _{”}
24.78644 1 20 09.06 +46 55 47.5  Šarounová
24.78924 1 20 08.37 +46 55 55.4  _{”}
24.79205 1 20 07.67 +46 56 03.1  _{”}
24.79428 1 20 07.15 +46 56 09.0  Korlević
24.79486 1 20 06.97 +46 56 10.8 13.5 Šarounová
24.79786 1 20 06.29 +46 56 19.2  Korlević
24.80102 1 20 05.45 +46 56 28.1  _{”}
24.80641 1 20 04.13 +46 56 43.0  Tichá
24.80759 1 20 03.84 +46 56 46.2  _{”}
24.80863 1 20 03.58 +46 56 49.1 13.4  _{”}
24.80984 1 20 03.28 +46 56 52.5  _{”}

M. Bezpalko, M. Blythe, F. Shelly, and M. Elowitz (Lincoln Laboratory ETS, Socorro). Measurers J. Stuart, H. Viggh, and R. Sayer. 1.0-m f/2.15 reflector + CCD. Tail in p.a. 200^{o}; magnitude 15.6–15.7.

L. Šarounová, D. Ondřich, and M. Prouza (Ondřejov). 0.65-m f/3.6 reflector + CCD. Coma diameter 35′′; rather wide tail 3′ long in p.a. 200^{o}; m_{2} = 14.8.

K. Korlević (Višnjan). 0.41-m f/4.3 reflector + CCD. Coma diameter 40′′;

tail 60′′ long in p.a. 160^{o}–190^{o}; magnitude 14.8. J. Tichá and M. Tichý (Kleť). 0.57-m f/5.2 reflector + CCD. Tail > 18′′

long in p.a. 225^{o}.

SUPERNOVA 1999dq IN NGC 976

CCD magnitude estimates, unfiltered unless otherwise noted: Sept. 4.719 UT, 15.2 (S. Yoshida and K. Kadota, Ageo, Saitama, Japan); 10.130, V = 14.28 ± 0.02 (D. Hanžl, Brno, Czech Republic); 11.076, 15.2 (E. Prosperi, Larciano, Italy).

1999 September 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green